# Berms without Calibration

### v1

@@ 2 Swaptions with Relative Strikes
When a vanilla swaption is traded, its strike is typically quoted as a spread relative to the prevailing ATM level - the par rate of the underlying swap. Once the swaption contract is agreed and entered by the counterparties the product is booked with an absolute strike equal to the sum of ATM level and the spread. In this section we consider a modification to the swaption contract where the ATM level is also a part of the contract and its fixing time $T_{fix}$ can be any time between the contract issue date and the swaption exercise date $T_{ex}$.
Consider first the case when the ATM level - the par rate of the underlying swap - is fixed at expiry $T_{ex}$, i.e. the absolute strike of the swaption becomes only known/fixed at the swaption expiry. This payoff is deterministic
@@
The expectation above is with respect to the joint distribution of the swap rate $R(T_{s},T_{s},T_{e})$ as fixed at $T_{s}$ and the swap rate $R(T_{fix},T_{s},T_{e})$ as fixed at $T_{fix}$ and can be performed by copula integration. The probability density for the swap rate $R(T_{s},T_{s},T_{e})$ as fixed at $T_{s}$ is the standard distribution observed in the vanilla swaption market. The distribution of the swap rate $R(T_{fix},T_{s},T_{e})$ as fixed at $T_{fix}$ is the midcurve swaption rate distribution modelled as a joint distribution of $R(T_{fix},T_{fix},T_{e})$ as fixed at $T_{fix}$ and $R(T_{fix},T_{fix},T_{s})$ as fixed at $T_{fix}$, both are the standard swap rate distributions observed in the vanilla swaption market. To use the copula integration, one would need to estimate the correlation
@@
Under Gaussian assumptions the fair value of a swaption with relative strike $K$ is:
@@
As there is no developed market for the relative strike swaptions the adequate management of the product requires a dedicated parameter. To preserve the relation with midcurves the parameter can be managed as a multiplicative spread on top of the relative strike swaption volatility $\sqrt{\sigma^{2}_{x}-\sigma^{2}_{z}}$. To not burden the explanation, we shall not use the parameter in the following sections, as all the derivations can be easily adjusted.
@@ 3 Canary Swaptions
$\frac{C(t,T_{1},T_{2},T_{e},K)}{A(t,T_{2},T_{e})}$ $=$ $\mathbb{E}^{{\cal A}_{2}}\Bigl[max\Bigl\{\frac{A(T_{1},T_{1},T_{e})}{A(T_{1},T_{2},T_{e})}\Bigl(\omega(R(T_{1},T_{1},T_{e})-K)\Bigr)^{+},$ (27)
$\mathbb{E}^{{\cal A}_{2}}[(\omega(R(T_{2},T_{2},T_{e})-K))^{+}|{\cal F}_{T_{1}}]\Bigr\}\Big|{\cal F}_{t}\Bigr].$
To evaluate the overall expectation the stochastic variable $(\omega(R(T_{1},T_{1},T_{e})-K))^{+}$ needs to be moved from its natural ${\cal A}_{1}$ annuity measure which corresponds to the annuity $A(T,T_{1},T_{e})$ of the first coterminal swap $S(T_{1},T_{e},K)$, to ${\cal A}_{2}$ annuity measure corresponding to the annuity $A(T,T_{2},T_{e})$ of the second cotermianl swap. The change of measure is the forward starting swaption $W(T_{1},S(T_{2},T_{e},K))$ measure change constructed in Section 1. In this section we shall assume that the annuity ratio $\frac{A(T_{1},T_{1},T_{e})}{A(T_{1},T_{2},T_{e})}$ is deterministic, but as discussed in the earlier section the general case can be treated using an appropriate convexity adjustment.
For the second expectation, i.e. the one inside the curly brackets corresponding to $R(T_{2},T_{2},T_{e})$ we can write:
@@
where $\sigma_{X_{2}}$ is the implied volatility at strike (inclusive the square root time factor) of the distribution $pdf_{X_{2}}$ for $X_{2}=R(T_{2},T_{2},T_{e})$ as observed at $T_{2}$ and $\sigma_{Z}$ is the implied volatility at strike (inclusive the square root time factor) of the distribution $pdf_{Z}$ for $Z=R(T_{1},T_{2},T_{e})$ as observed at $T_{1}$, $N(\mu,\sigma)$ is used to denote the relevant normal random variable with the mean $\mu$ and the variance $\sigma^{2}$, $\Omega(\cdot)$ is the Bachelier option price as described in Section 2. Note that in (LABEL:SecExp) we are using references to the normal distribution and Bachelier option formula for the convenience of the explanation. In general (LABEL:SecExp) is the option part of the payoff of a swaption with the relative strike as described in Section 2. One can use a multiplicative spread over $\sqrt{\sigma^{2}_{X_{2}}-\sigma^{2}_{Z}}$ or a copula integration with the correlation from Theorem 3 to express the dependence on the volatility smile.
@@
$\frac{C(t,T_{1},T_{2},T_{e},K)}{A(t,T_{2},T_{e})}$ $=$ $\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}max\Bigl\{\frac{A(t,T_{1},T_{e})}{A(t,T_{2},T_{e})}\left(\omega(x-K)\right)^{+},\Omega\left(0,K-z,\sqrt{\sigma^{2}_{X_{2}}-\sigma^{2}_{Z}},\omega\right)\Bigr\}\times$ (29)
@@
so that in the deterministic annuities ratios case the volatility of $Z=R(T_{1},T_{2},T_{e})$ is
@@
where $\rho_{2,e}$ is the midcurve correlation (between $R(T_{1},T_{1},T_{e})$ and $R(T_{1},T_{1},T_{2})$).
@@
and the option (LABEL:SecExp) becomes deterministic, i.e. is the intrinsic value. In this case the Canary swaption payoff (27) is approximated by the maximum of two correlated normal variables:
@@
[…]
We performed a comparison of valuations based on Theorem 5 and Corollary 1 in real market conditions observed close of 30th June 2025 against the classical Hagan LGM. The volatility surface used for all three models is SABR surface calibrated to the observed market. The valuation based on Theorem 5 is performed using Simpson rule based double integration, the maximum in Corollary 1 is evaluated by the moment matching techniques as outlined in [10]. Figure 1 shows the difference of each of three valuations with the market consensus for a range of available strikes between ATM-200bps and ATM+600bps for a payer callable swap which starts in 2 years and can be exercised annually into either 2 or 1 year swap. This is the product equivalent to a Canary swaption 2Y1Y2Y described in this section. The doted lines are plus/minus one standard deviations of the market quotes.
Figure 1: Model price vs market consensus under different model assumptions: the callable swap equivalent to the Canary swaption on a 2Y swap starting in 2Y with the annual exercise frequency. Long/short swap rates correlation 89%.
@@ 4 Berms
The result is a stochastic variable over a joint distribution of $\mathbb{E}^{{{\cal A}_{n}}}[R(T_{i+k},T_{i+k},T_{e})|{\cal F}_{T_{i}}]$, k=0,nik=0,...n-i. The distribution of the random variable $\mathbb{E}^{{\cal A}_{n}}[R(T_{i+k},T_{i+k},T_{e})|{\cal F}_{T_{i}}]$ is the underlying (forward starting) swap rate distribution for a midcurve swaption on $R(T_{i},T_{i+k},T_{e})$ but observed in (unnatural) annuity measure ${\cal A}_{n}$}. The measure change is trivial under the deterministic annuity ratio assumption while in the general case it can be handled with an appropriate convexity adjustment. To find the correlation matrix of this joint distribution we can use an alternative derivation for the Canary swaption payoff expectation as an integral over a joint probability density of various swap rates as follows. Consider for any $i,j$ satisfying $1\leq i<j\leq n$, the swap rate distributions
@@
When integrating the swaption payoff along the joint distribution, the distribution $pdf_{Y}$ corresponding to $R(T_{i},T_{i},T_{j})$ will be integrated out as the payoff does not explicitly depend on it. Nevertheless, the impact of $R(T_{i},T_{i},T_{j})$ will still be felt through the correlation. There are two sources of the correlation in this model. The correlation $\rho_{X_{i},Y}$ between $R(T_{i},T_{i},T_{e})$ and $R(T_{i},T_{i},T_{j})$ which describes $R(T_{i},T_{j},T_{e})$ as a joint density of $pdf_{X_{i}}$ and $pdf_{Y}$ and the correlation $\rho_{X_{j}}$ between $R(T_{j},T_{j},T_{e})$ observed at $T_{j}$ with $R(T_{i},T_{j},T_{e})$ observed at $T_{i}$. The latter coincides with $\rho(T_{i},T_{j},R(T_{j},T_{e}))$ as defined in Theorem 3. We can write the relationship in terms of independent Brownian motions $W^{X_{i}}_{T_{i}}$, $W^{Y}_{T_{i}}$, $W^{Z}_{T_{i}}$, $W^{X_{j}}_{T_{i}}$, and $W^{X_{j}}_{T_{i}T_{j}}$:
$X_{i}$ $=$ $\tilde{\sigma}_{X_{i}}(\sqrt{1-\rho^{2}_{X_{i},Y}}W^{X_{i}}_{T_{i}}+\rho_{X_{i},Y}W^{Y}_{T_{i}}),$
$Y$ $=$ $\tilde{\sigma}_{Y}W^{Y}_{T_{i}},$
$Z$ $=$ $\tilde{\sigma}_{Z}W^{Z}_{T_{i}},$
$X_{j}$ $=$ $\tilde{\sigma}_{X_{j}}\left(W^{X_{j}}_{T_{i}T_{j}}+\rho_{X_{j}}W^{Z}_{T_{i}}+\sqrt{1-\rho^{2}_{X_{j}}}W^{X_{j}}_{T_{i}}\right),$ (40)
where by $\tilde{\sigma}$ we underline that the volatilities do not include the square root time factor. Equating
@@
Under the assumption of the deterministic ratios of annuities and perfect correlations in all the forward rates, the Bermudan swaption payoff is
@@
Proof: If we work under the assumption of deterministic ratios of annuities, then we can replace each $\frac{A(T_{i},T_{i},T_{e})}{A(T_{i},T_{n},T_{e})}$ in (37) by $\frac{A(t,T_{i},T_{e})}{A(t,T_{n},T_{e})}$, use pdfs of the individual swap rates in their own annuity measures and evaluate the payoff as (LABEL:BPC). For the correlation matrix we can simplify (42) because $\rho_{X_{j}}=1$ and for any $1\leq i<j\leq n$
@@
$\times$ $\frac{A(t,T_{i},T_{e})^{2}\sigma^{2}_{X_{i}}+A(t,T_{i},T_{j})^{2}\sigma^{2}_{R(T_{i},T_{i},T_{j})}-A(t,T_{j},T_{e})^{2}\sigma^{2}_{X_{j}}}{2A(t,T_{i},T_{e})A(t,T_{i},T_{j})\sigma_{X_{i}}\sigma_{R(T_{i},T_{i},T_{j})}}$
@@
Note that as in the Corollary 1 if we assume that all the distributions are Gaussian then the integration can be done by the moment matching techniques from [10].
@@
as a function $f^{(j)}(\cdot)$ of $x_{j}=\mathbb{E}^{{\cal A}_{j}}[R(T_{j},T_{j},T_{e})|{\cal F}_{T_{j}}]$, where ${\cal A}_{j}$ is the annuity meassure corresponding to the $j$th coterminal swap $S(T_{j},T_{e},K)$ and by the superscript $(j)$ we express that the only part of the swaption active is the one to be exercised in to coterminals starting from time $T_{j}$ and beyond.
@@
We performed a comparison of valuations based on Theorem 7 and Theorem 8 in real market conditions observed close of 30th June 2025 against the classical Hagan LGM. The volatility surface used for all three models is SABR surface calibrated to the observed market. The maximum in Theorem 7 is evaluated by the moment matching techniques as outlined in [10]. A direct application of the Theorem 7 results in valuations which are significantly below the market. To demonstrate the qualitative behaviour of the corresponding model we adjusted the correlation close to 100%. The valuation based on Theorem 8 is performed using a) Hagan lattice step for the last inside expectation $\mathbb{E}^{{{}_{A}}_{j}}\Bigl[f^{(j)}(z+\tilde{x}_{j})|{\cal F}_{T_{j-1}}\Bigr]$ and b) Simpson rule integration for the overall conditional expectation (the middle expectation is used to express the functional dependence on two variables: $x_{j-1}$ and $\tilde{x}_{j}=\mathbb{E}^{{{}_{A}}_{j}}[R(T_{j},T_{j},T_{e})|{\cal F}_{T_{j-1}}]$). The overall conditional expectation is a function of $x_{j-1}$ and is re-used at the next step of Hagan lattice scheme. Figure 2 shows the difference of each of three valuations with the market consensus for a range of strikes between ATM-300bps and ATM+600bps available for a payer callable swap which starts in 5 years, last 10 years and has the annual exercise. The doted lines are plus/minus one standard deviations of the individual market quotes.
Figure 2: Model price vs market consensus under different model assumptions: the callable swap equivalent to the Bermudan swaption on 10Y swap starting in 5Y with the annual exercise frequency . Linear coterminals vs 1Y swap rate correlation term structure from 88% to 99%. Moments matching based on 99.9% correlations.

### v2

@@ 2 Swaptions with Relative Strikes
When a vanilla swaption is traded, its strike is typically quoted as a spread relative to the prevailing ATM level - the par rate of the underlying swap. Once the swaption contract is agreed and entered by the counterparties, the product is booked with an absolute strike equal to the sum of ATM level and the spread. In this section we consider a modification to the swaption contract where the ATM level is also a part of the contract and its fixing time $T_{fix}$ can be any time between the contract issue date and the swaption exercise date $T_{ex}$.
First, consider the case when the ATM level - the par rate of the underlying swap - is fixed at expiry $T_{ex}$, i.e. the absolute strike of the swaption becomes only known/fixed at the swaption expiry. This payoff is deterministic
@@
The expectation above is with respect to the joint distribution of the swap rate $R(T_{s},T_{s},T_{e})$ as fixed at $T_{s}$ and the swap rate $R(T_{fix},T_{s},T_{e})$ as fixed at $T_{fix}$ and can be performed by copula integration. The probability density for the swap rate $R(T_{s},T_{s},T_{e})$ as fixed at $T_{s}$ is the standard distribution observed in the vanilla swaption market. The distribution of the swap rate $R(T_{fix},T_{s},T_{e})$ as fixed at $T_{fix}$ is the midcurve swaption rate distribution modelled as a joint distribution of $R(T_{fix},T_{fix},T_{e})$ as fixed at $T_{fix}$ and $R(T_{fix},T_{fix},T_{s})$ as fixed at $T_{fix}$, both are the standard swap rate distributions observed in the vanilla swaption market. To use the copula integration, one will need to estimate the correlation
@@
Under Gaussian assumptions the fair value of a swaption with a relative strike $K$ is:
@@
As there is no developed market for the relative strike swaptions, the adequate management of the product requires a dedicated parameter. To preserve the relation with midcurves, the parameter can be managed as a multiplicative spread on top of the relative strike swaption volatility $\sqrt{\sigma^{2}_{x}-\sigma^{2}_{z}}$. To not burden the explanation, we shall not use the parameter in the following sections, as all the derivations can be easily adjusted.
@@ 3 Canary Swaptions
$\frac{C(t,T_{1},T_{2},T_{e},K)}{A(t,T_{2},T_{e})}=$
$\mathbb{E}^{{\cal A}_{2}}\Bigl[max\Bigl\{\frac{A(T_{1},T_{1},T_{e})}{A(T_{1},T_{2},T_{e})}\Bigl(\omega(R(T_{1},T_{1},T_{e})-K)\Bigr)^{+},\mathbb{E}^{{\cal A}_{2}}\Bigl[\Bigl(\omega(R(T_{2},T_{2},T_{e})-K)\Bigr)^{+}|{\cal F}_{T_{1}}\Bigr]\Bigr\}\Big|{\cal F}_{t}\Bigr].$
To evaluate the overall expectation the stochastic variable $(\omega(R(T_{1},T_{1},T_{e})-K))^{+}$ needs to be moved from its natural ${\cal A}_{1}$ annuity measure, which corresponds to the annuity $A(T,T_{1},T_{e})$ of the first coterminal swap $S(T_{1},T_{e},K)$, to ${\cal A}_{2}$ annuity measure corresponding to the annuity $A(T,T_{2},T_{e})$ of the second cotermianl swap. The change of measure is the forward starting swaption $W(T_{1},S(T_{2},T_{e},K))$ measure change constructed in Section 1. In this section we will assume that the annuity ratio $\frac{A(T_{1},T_{1},T_{e})}{A(T_{1},T_{2},T_{e})}$ is deterministic. In the general case the Canary swaption payoff (30) can be treated within Theorem 2 as the payoff formula is a function of $\frac{A(T_{1},T_{1},T_{e})}{A(T_{1},T_{2},T_{e})}(R(T_{1},T_{1},T_{e})-K)$ and $\frac{A(T_{1},T_{1},T_{2})}{A(T_{1},T_{2},T_{e})}(R(T_{1},T_{1},T_{2})-K)$.
For the second expectation in (30), i.e. the one inside the curly brackets and corresponding to $R(T_{2},T_{2},T_{e})$, we can write:
@@
where $\sigma_{X_{2}}$ is the implied volatility at strike (inclusive the square root time factor) of the distribution $pdf_{X_{2}}$ for $X_{2}=R(T_{2},T_{2},T_{e})$ (i.e. as observed at $T_{2}$) and $\sigma_{Z}$ is the implied volatility at strike (inclusive the square root time factor) of the distribution $pdf_{Z}$ for $Z=R(T_{1},T_{2},T_{e})$ (i.e. as observed at $T_{1}$); $N(\mu,\sigma)$ is used to denote the relevant normal random variable with the mean $\mu$ and the variance $\sigma^{2}$, $\Omega(\cdot)$ is the Bachelier option price as described in Section 2. Note that in (LABEL:SecExp) we are using references to the normal distribution and Bachelier option formula for the convenience of the explanation. In general (LABEL:SecExp) is the option part of the payoff of a swaption with the relative strike as described in Section 2. One can use a multiplicative spread over $\sqrt{\sigma^{2}_{X_{2}}-\sigma^{2}_{Z}}$ or a copula integration with the correlation from Theorem 3 to express the dependence on the volatility smile.
@@
$\frac{C(t,T_{1},T_{2},T_{e},K)}{A(t,T_{2},T_{e})}$ $=$ $\int^{\infty}_{-\infty}\int^{\infty}_{-\infty}max\Bigl\{\frac{A(t,T_{1},T_{e})}{A(t,T_{2},T_{e})}\left(\omega(x-K)\right)^{+},\Omega\left(0,K-z,\sqrt{\sigma^{2}_{X_{2}}-\sigma^{2}_{Z}},\omega\right)\Bigr\}$ (32)
@@
so that for the deterministic annuity ratios, the volatility of $Z=R(T_{1},T_{2},T_{e})$ is
@@
where $\rho_{2,e}$ is the midcurve correlation (i.e. between $R(T_{1},T_{1},T_{e})$ and $R(T_{1},T_{1},T_{2})$).
@@
and the option (LABEL:SecExp) becomes deterministic, i.e. the intrinsic value. In this case the Canary swaption payoff (30) is approximated by the maximum of the following two correlated normal variables:
@@
In the next sections we will refer to the condition of the Corollary 1 as the condition of perfect correlations of the forward rates. This is because $Z$ and $X_{2}$ are distributions of the same swap rate but observed at different times in the future.
@@
We performed a comparison of valuations based on Theorem 5 and Corollary 1 in real market conditions observed close of business day on 30 June 2025 against the classical Hagan LGM. The volatility surface used for all three models is SABR surface calibrated to the observed European swaption market. The valuation based on Theorem 5 is performed using Simpson’s rule for the double integral; the maximum in Corollary 1 is evaluated by the moment matching techniques as outlined in [11]. The correlation between the long and the short swap rates is 89%. To account for the impact of the volatility smile we chose volatilities at relevant strikes when evaluating individual Bermudan prices via (32) or (LABEL:MntCF). Figure 1 shows the difference of each of the three valuations with the market consensus for a range of available strikes between ATM-200bps and ATM+600bps for a payer callable swap, which starts in 2 years and can be exercised annually into either 2 or 1 year swap. This is the product equivalent to a Canary swaption 2Y1Y2Y described in this section. The doted lines are plus/minus one standard deviations of the market quotes.
Figure 1: Model price vs market consensus under different model assumptions: the callable swap equivalent to the Canary swaption on a 2Y swap starting in 2Y with the annual exercise frequency.
@@ 4 Berms
The result is a stochastic variable over a joint distribution of $\mathbb{E}^{{{\cal A}_{n}}}[R(T_{i+k},T_{i+k},T_{e})|{\cal F}_{T_{i}}]$, k=0,nik=0,...n-i. The distribution of the random variable $\mathbb{E}^{{\cal A}_{n}}[R(T_{i+k},T_{i+k},T_{e})|{\cal F}_{T_{i}}]$ is the underlying (forward starting) swap rate distribution for a midcurve swaption on $R(T_{i},T_{i+k},T_{e})$ but observed in (unnatural) annuity measure ${\cal A}_{n}$}. The measure change is trivial under the deterministic annuity ratio assumption while in the general case it can be handled with an appropriate convexity adjustment using Theorem 2 . To find the correlation matrix of this joint distribution we can use an alternative derivation for the Canary swaption payoff expectation as an integral over a joint probability density of various swap rates as follows. Consider for any $i,j$ satisfying $1\leq i<j\leq n$, the swap rate distributions
@@
When integrating the swaption payoff along the joint distribution, the distribution $pdf_{Y}$ corresponding to $R(T_{i},T_{i},T_{j})$ will be integrated out as the payoff does not explicitly depend on it. Nevertheless, the impact of $R(T_{i},T_{i},T_{j})$ will still be felt through the correlation. There are two sources of the correlation in this model. The correlation $\rho_{X_{i},Y}$ between $R(T_{i},T_{i},T_{e})$ and $R(T_{i},T_{i},T_{j})$, which describes $R(T_{i},T_{j},T_{e})$ as a joint density of $pdf_{X_{i}}$ and $pdf_{Y}$, and the correlation $\rho_{X_{j}}$ between $R(T_{j},T_{j},T_{e})$ observed at $T_{j}$ with $R(T_{i},T_{j},T_{e})$ observed at $T_{i}$. The latter coincides with $\rho(T_{i},T_{j},R(T_{j},T_{e}))$ as defined in Theorem 3. We can write the relationship in terms of independent Brownian motions $W^{X_{i}}_{T_{i}}$, $W^{Y}_{T_{i}}$, $W^{Z}_{T_{i}}$, $W^{X_{j}}_{T_{i}}$, and $W^{X_{j}}_{T_{i}T_{j}}$:
$X_{i}$ $=$ $(\cdots)+\tilde{\sigma}_{X_{i}}(\sqrt{1-\rho^{2}_{X_{i},Y}}W^{X_{i}}_{T_{i}}+\rho_{X_{i},Y}W^{Y}_{T_{i}}),$
$Y$ $=$ $(\cdots)+\tilde{\sigma}_{Y}W^{Y}_{T_{i}},$
$Z$ $=$ $(\cdots)+\tilde{\sigma}_{Z}W^{Z}_{T_{i}},$
$X_{j}$ $=$ $(\cdots)+\tilde{\sigma}_{X_{j}}\left(W^{X_{j}}_{T_{i}T_{j}}+\rho_{X_{j}}W^{Z}_{T_{i}}+\sqrt{1-\rho^{2}_{X_{j}}}W^{X_{j}}_{T_{i}}\right),$ (43)
where by $\tilde{\sigma}$ we underline that the volatilities do not include the square root time factor and by $(\cdots)$ we suppress the drifts (which do not impact correlations). Equating (with omitted drifts)
@@
Under the assumption of the deterministic ratios of annuities and perfect correlations in all the forward rates, Bermudan swaption payoff is
@@
Proof: If we work under the assumption of deterministic ratios of annuities, then we can replace each $\frac{A(T_{i},T_{i},T_{e})}{A(T_{i},T_{n},T_{e})}$ in (40) by $\frac{A(t,T_{i},T_{e})}{A(t,T_{n},T_{e})}$, use pdfs of the individual swap rates in their own annuity measures and evaluate the payoff as (LABEL:BPC). For the correlation matrix we can simplify (45) because $\rho_{X_{j}}=1$, and for any $1\leq i<j\leq n$
@@
$\frac{A(t,T_{i},T_{e})^{2}\sigma^{2}_{X_{i}}+A(t,T_{i},T_{j})^{2}\sigma^{2}_{R(T_{i},T_{i},T_{j})}-A(t,T_{j},T_{e})^{2}\sigma^{2}_{X_{j}}}{2A(t,T_{i},T_{e})A(t,T_{i},T_{j})\sigma_{X_{i}}\sigma_{R(T_{i},T_{i},T_{j})}}$
@@
as a function $f^{(j)}(\cdot)$ of $x_{j}=\mathbb{E}^{{\cal A}_{j}}[R(T_{j},T_{j},T_{e})|{\cal F}_{T_{j}}]$, where ${\cal A}_{j}$ is the annuity measure corresponding to the $j$th coterminal swap $S(T_{j},T_{e},K)$ and by the superscript $(j)$ we express that the only part of the swaption active is the one to be exercised in to coterminals starting from time $T_{j}$ and beyond.
@@
We performed a comparison of valuations based on Theorem 7 and Theorem 8 in real market conditions observed close of business day on 30 June 2025 against the classical Hagan LGM. The volatility surface used for all three models is SABR surface calibrated to the observed European swaption market. The maximum in Theorem 7 is evaluated by the moment matching techniques as outlined in [11]. A direct application of Theorem 7 results in valuations which are significantly below the market. To demonstrate the qualitative behaviour of the corresponding model we adjusted the correlation close to 100%. The valuation based on Theorem 8 is performed using a) Hagan lattice step for the last inside expectation $\mathbb{E}^{{{}_{A}}_{j}}\Bigl[f^{(j)}(z+\tilde{x}_{j})|{\cal F}_{T_{j-1}}\Bigr]$ and b) Simpson rule integration for the overall conditional expectation (the middle expectation is used to express the functional dependence on two variables: $x_{j-1}$ and $\tilde{x}_{j}=\mathbb{E}^{{{}_{A}}_{j}}[R(T_{j},T_{j},T_{e})|{\cal F}_{T_{j-1}}]$). The overall conditional expectation is a function of $x_{j-1}$ and is re-used at the next step of Hagan lattice scheme. We use a linear term structure ranging from 88% to 99% for the correlations between the coterminal swap rates and 1Y swap rates. The moments matching is based on 99.9% correlations. Figure 2 shows the difference of each of three valuations with the market consensus for a range of strikes between ATM-300bps and ATM+600bps available for a payer callable swap which starts in 5 years, last 10 years and has the annual exercise. The doted lines are plus/minus one standard deviations of the individual market quotes.
Figure 2: Model price vs market consensus under different model assumptions: the callable swap equivalent to the Bermudan swaption on 10Y swap starting in 5Y with the annual exercise frequency.